i need help solving linear equations. How do i solve -1/2x+5(x-3)=x-1

Answer by geetha_rama(94)

-1/2x + 5(x-3) = x-1
5(x-3)=x+x/2 -1
5x - 15 + 1 = (2x+x)/2
5x - 14 = 3x/2
10x - 28 = 3x
=> 7x = 28
x = 4